LALL IU HAL BRIR PRIRTEA VERION BACK Problem 8-24 (Part Level Submission) Wildhorse Water Co. is a leading producer of greenh thank you!
Answer #1

Total relevant cost (Make) = ($12+7+4+9*0.30)*40960 = $1052672

Total relevant cost (Buy) = 36*40960 = $1474560